### 17 Mar 2026 v1

# OPERA: Online Data Pruning for Efficient Retrieval Model Adaptation

Abstract
Domain-specific finetuning is essential for dense retrievers, yet not all training pairs contribute equally to the learning process. We introduce OPERA, a data pruning framework that exploits this heterogeneity to improve both the effectiveness and efficiency of retrieval model adaptation. We first investigate static pruning (SP), which retains only high-similarity query-document pairs, revealing an intrinsic quality-coverage tradeoff: ranking (NDCG) improves while retrieval (Recall) can degrade due to reduced query diversity. To resolve this tradeoff, we propose a two-stage dynamic pruning (DP) strategy that adaptively modulates sampling probabilities at both query and document levels throughout training, prioritizing high-quality examples while maintaining access to the full training set. Evaluations across eight datasets spanning six domains demonstrate the effectiveness of both approaches: SP improves ranking over standard finetuning (NDCG@10 +0.5%), while DP achieves the strongest performance on both ranking (NDCG@10 +1.9%) and retrieval (Recall@20 +0.7%), with an average rank of 1.38 across all methods. These findings scale to Qwen3-Embedding, an LLM-based dense retriever, confirming architecture-agnostic benefits. Notably, DP reaches comparable performance in less than 50% of the training time required by standard finetuning.
1 Introduction
Dense retrievers have advanced information retrieval (Izacard et al., 2021; Reimers and Gurevych, 2019; Karpukhin et al., 2020; Xiao et al., 2023; Muennighoff et al., 2024; Wang et al., 2022; Lee et al., 2024), substantially outperforming traditional sparse methods (Robertson et al., 2009; Ramos and others, 2003). Built on pretrained language models (Vaswani et al., 2017; Devlin et al., 2018; Jiang et al., 2023), these models achieve strong zero-shot performance across diverse benchmarks (Thakur et al., 2021; Muennighoff et al., 2022). Nevertheless, achieving optimal performance on specific downstream tasks still requires domain-specific finetuning (Thakur et al., 2022; Howard and Ruder, 2018).
Data pruning and coreset selection have shown promise for improving training efficiency in neural networks (Toneva et al., 2018; Marion et al., 2023; Sorscher et al., 2022; Killamsetty et al., 2021), and dynamic pruning methods have further demonstrated that adjusting data selection during training can maintain performance while reducing computation (Qin et al., 2023; Li et al., 2024; Huang et al., 2024). However, these methods are designed for standard classification or generation tasks where training samples are treated as independent, identically distributed instances. Dense retriever finetuning is fundamentally different: it employs a two-stage contrastive sampling framework (Xiao et al., 2023; Chen et al., 2024; Zhang et al., 2023) where queries are first sampled, and then positive and negative documents are selected for each query. This hierarchical structure means that data quality operates at two distinct granularities (query relevance and document relevance), creating unique challenges that existing pruning methods do not address. To our knowledge, no prior work has studied data pruning specifically for dense retriever finetuning. A detailed discussion of related work is provided in Appendix A.
We introduce OPERA, a framework that exploits the heterogeneous quality of training data (Tirumala et al., 2024; Abbas et al., 2023) to improve domain adaptation for dense retrievers. Our investigation begins with static pruning (SP), which retains only the highest-similarity query-document pairs for training. This simple strategy reveals a key insight: quality-based filtering consistently improves ranking metrics (NDCG) but can degrade retrieval coverage (Recall), because pruning disproportionately removes queries with fewer high-quality documents, breaking the balanced sampling that retrievers rely on for broad coverage. This quality-coverage tradeoff is intrinsic to the two-stage sampling structure of retrieval training and motivates the need for a more nuanced approach.
To resolve this tradeoff, we propose dynamic pruning (DP), which maintains the complete training set while adaptively adjusting sampling probabilities at both query and document levels. Unlike InfoBatch (Qin et al., 2023), which uses a fixed loss-average threshold and rescales gradients to maintain unbiasedness, our approach implements dynamic thresholds that evolve throughout training via cosine scheduling, and preserves original learning rates to emphasize high-quality training signals. Rather than employing hard exclusions, DP assigns reduced but nonzero sampling probabilities to lower-quality examples, ensuring continued data diversity while prioritizing informative instances.
Concretely, our dynamic pruning framework combines three components: (1) hierarchical pruning at both query and document granularities, reflecting the two-stage sampling structure; (2) dynamic threshold scheduling that progressively sharpens selection as model representations improve; and (3) soft pruning mechanisms that modulate sampling probabilities rather than discarding data.
We evaluate OPERA on eight datasets spanning nutrition (Boteva et al., 2016; Thakur et al., 2021), medicine (Rekabsaz et al., 2021), finance (Maia et al., 2018; Thakur et al., 2021), non-factoid QA (Hashemi et al., 2020), factoid QA (Joshi et al., 2017; Karpukhin et al., 2020; Yang et al., 2018; Thakur et al., 2021), and fact verification (Thorne et al., 2018; Thakur et al., 2021), including both datasets seen and unseen during pretraining. Our contributions are:
  • We identify a quality-coverage tradeoff unique to retrieval’s two-stage sampling: static quality filtering improves ranking but degrades recall. This finding holds across encoder-only (BGE) and LLM-based (Qwen3-Embedding) retrievers.
  • We propose dynamic pruning with hierarchical scheduling that resolves this tradeoff, improving both ranking and recall while halving convergence time. We provide an efficient formulation compatible with fixed-iteration training frameworks.
  • We validate OPERA across 8 datasets, 6 domains, and 2 architectures (encoder-only BGE and decoder-based Qwen3-Embedding), with theoretical guarantees on when pruning outperforms standard finetuning.
2 Methodology
2.1 Preliminary: Standard Finetuning (FT)
Given a dataset $Q$ with $n$ queries where each query $q\in Q$ has $m_{q}$ positive documents, we adopt a contrastive learning framework with two-stage sampling (Xiao et al., 2023). For each training step, we first uniformly sample queries from $Q$. Then, for each sampled query, we randomly select one positive document $d$ from its $m_{q}$ positive documents and one negative document through hard negative mining for loss computation.
The sampling probabilities are:
$P_{t}(q)=\nicefrac{{1}}{{n}},\quad P_{t}(d|q)=\nicefrac{{1}}{{m_{q}}}$ (1)
Note that for notational simplicity, we omit the time step $t$ for the following equations.
2.2 Static Pruning (SP)
To understand how data quality affects retrieval finetuning, we begin with a straightforward approach inspired by Sorscher et al. (2022): retaining only the highest-quality training pairs. We compute cosine similarity between each query-document pair using the pretrained model and retain only the top fraction of pairs for training. Mathematically, we use an indicator function $I$ to mark whether a query or query-document pair is kept:
$I=\begin{cases}1,&\text{if kept}\\ 0,&\text{if pruned}\end{cases}$ (2)
This adjusts our sampling probabilities to:
$P(q)=\frac{\sum_{d}I_{d}}{\sum_{q}\sum_{d}I_{d}},\quad P(d|q)=\frac{I_{d}}{\sum I_{d}}$ (3)
In practice, we select top $k\sum m_{q}$ query-document pairs with highest similarity scores for training, where $k$ is the data retention rate:
$k=\frac{\sum_{q}\sum_{d}I_{d}}{\sum m_{q}}$ (4)
Crucially, quality-based filtering breaks the uniform query sampling that standard finetuning relies on: queries with more high-similarity documents are overrepresented, while queries with fewer such documents may be excluded entirely. This improves ranking (e.g., NDCG) by focusing on well-matched pairs, but can degrade recall by reducing coverage of the query space. This quality-coverage tradeoff is intrinsic to the two-stage sampling structure of retrieval training and motivates the dynamic approach described in Section 2.3.
Alternative scoring metric. We also evaluate consistency-based scoring (CBS) (Wang et al., 2022), which ranks each positive pair against random negatives and retains pairs that consistently rank highly, adapted to our finetuning setting. CBS and cosine similarity yield comparable results (Appendix D.1), but CBS is incompatible with DP as it requires re-embedding all negative documents after each model update. We therefore adopt cosine similarity as the default metric for both SP and DP.
2.3 Dynamic Pruning (DP)
Algorithm 1 Dynamic Pruning
procedure OPERA
  $\alpha_{s}\leftarrow\text{query strength start}$
  $\alpha_{e}\leftarrow\text{query strength end}$
  $r_{s}\leftarrow\text{query ratio start}$
  $\beta_{s}\leftarrow\text{doc strength start}$
  $\beta_{e}\leftarrow\text{doc strength end}$
  $v_{s}\leftarrow\text{doc ratio start}$
  $v_{e}\leftarrow\text{doc ratio end}$
  $n\leftarrow\text{size of dataset}$
  $t_{max}\leftarrow\text{max training steps}$
  $n_{0}=\lfloor n\cdot(1-r_{s})/\alpha_{s}+r_{s}\cdot n\rfloor$
  for each iteration do
   $t\leftarrow\text{current training step}$
   Update $\mathbf{q}$ $\triangleright$ Query scores
   Update $\mathbf{p}$ $\triangleright$ Pos. scores
   $\textsc{SampleQuery}()$
   $\textsc{SampleDocument}()$
  end for
end procedure
procedure SampleQuery
  $\alpha\leftarrow\alpha_{e}+\frac{(1+\cos(\frac{t}{t_{max}}\pi))\cdot(\alpha_{s}-\alpha_{e})}{2}$
  $r\leftarrow\frac{\alpha\cdot n_{0}-n}{(\alpha-1)\cdot n}$
  $q_{top}\leftarrow\text{Top $r$ Queries}$
  $q_{rem}\leftarrow queries\setminus q_{top}$
  $q_{rand}\leftarrow\text{Random Select $n_{0}-r$ from $q_{rem}$}$
  $q^{\prime}\leftarrow q_{top}\cup q_{rand}$
  Sample from $q^{\prime}$
end procedure
procedure SampleDocument
  $\beta\leftarrow\beta_{e}+\frac{(1+\cos(\frac{t}{t_{max}}\pi))\cdot(\beta_{s}-\beta_{e})}{2}$
  $v\leftarrow v_{e}+\frac{(1+\cos(\frac{t}{t_{max}}\pi))\cdot(v_{s}-v_{e})}{2}$
  Calculate the threshold $T$ at cutoff $v$
  $w=(\mathbf{p}>T)\cdot(\beta-1)+1$
  $w=normalize(w)$
  Sample documents with weights $w$
end procedure
The quality-coverage tradeoff identified in SP arises because hard pruning permanently discards data, reducing the diversity of training signals. Dynamic Pruning (DP) resolves this tradeoff by replacing hard exclusions with soft sampling modulation: high-quality examples receive elevated sampling probabilities, while lower-quality examples remain accessible with reduced frequencies. This preserves the broad data coverage needed for recall while concentrating training effort on informative instances.
The sampling process, detailed in Algorithm 1, consists of two components. For query sampling, we combine top-scoring queries with randomly selected low-scoring queries to maintain diversity. In document sampling, we assign sampling weights to documents based on their quality while ensuring all documents maintain a non-zero probability of selection.
The sampling probabilities for queries and query-document pairs are defined as:
[…]
Here, $n_{0}$ represents a fixed virtual dataset size, ensuring compatibility with training frameworks that require a static dataset size. $r_{q}$ is the fraction of high-quality positive pairs for each query. $\alpha$ and $\beta$ are sampling strength parameters that vary during training:
$\alpha(t)=\alpha_{e}+\frac{(1+\cos(\frac{t}{t_{max}}\pi))\cdot(\alpha_{s}-\alpha_{e})}{2}$ (6)
where $\alpha_{s}$ and $\alpha_{e}$ denote the start and end sampling strengths respectively, and $t_{max}$ is the maximum training steps. Similarly, $\beta$ follows the same cosine decay schedule.
Update Interval To mitigate the computational overhead associated with frequent score updates and pruning operations, we introduce an update interval parameter $I_{u}$. We update the query scores $\mathbf{q}$ every $I_{u}$ iterations, while maintaining the per-iteration random selection of $d$ due to its negligible computational cost. Our empirical analysis demonstrates that this optimization reduces additional computation time from 4.5% to 1.64% over the baseline, without impacting performance. The effects of varying $I_{u}$ are examined in Section 3.4.2.
2.4 Theoretical Analysis: When Does Pruning Help?
We formalize the conditions under which data pruning outperforms standard finetuning. In dense retrieval, a query $q$ and its positive document $d$ are encoded into normalized embeddings $e_{q},e_{d}\in\mathbb{R}^{h}$, with cosine similarity $s(e_{q},e_{d})=e_{q}^{T}e_{d}$. We analyze SP and DP by studying how each method’s sampling strategy affects the optimal query embedding when some positive labels are noisy. Throughout the analysis, we consider a single query $q$.
Lemma 1.
Let $u,v\in\mathbb{R}^{n}$ be unit vectors with $u\neq v$, and $k\in(0,1)$. Define $f(k)=\frac{(ku+(1-k)v)^{T}u}{\|ku+(1-k)v\|}.$ Then $f^{\prime}(k)>0$ for all $k\in(0,1)$.
Proof.
Direct computation yields $f^{\prime}(k)=\frac{(1-(u^{T}v)^{2})(1-k)}{\|ku+(1-k)v\|^{3}}.$ Since $u\neq v\implies u^{T}v<1$, and $k<1$, thus $f^{\prime}(k)>0$. ∎
Theorem 1.
Let $m_{q}$ be the total number of documents labeled as positive for query $q$, and $m_{q}^{+}$ be the number of correctly labeled documents ($m_{q}^{+}\leq m_{q}$). Assume correctly labeled documents have embedding mean direction $\mu_{1}\in\mathbb{R}^{h}$, noisy (false-positive) documents have mean direction $\mu_{2}\in\mathbb{R}^{h}$, and negative documents have mean $0\in\mathbb{R}^{h}$. For simplicity, we consider only the bias in the estimated query embedding and dismiss variance.
For SP, $r$ documents are selected ($\gamma r$ correctly labeled) for uniform sampling. For DP, $s$ documents are selected ($\rho s$ correctly labeled) with probability $\beta$ times higher than the remaining $m_{q}-s$ documents ($\beta>1$). Define $E^{\text{FT}}$, $E^{\text{SP}}$, $E^{\text{DP}}$ as the expected cosine similarity between the optimal query embedding and true-positive documents under each strategy.
Then:
$E^{SP}>E^{FT}\Leftrightarrow\gamma>\frac{m_{q}^{+}}{m_{q}}$ (7)
and similarly for DP. Additionally, when $\gamma=\rho$:
$E^{SP}>E^{DP}.$ (8)
Proof.
Under FT, the model maximizes $\mathcal{L}^{FT}=\sum_{i=1}^{m_{q}}s(e_{q},e_{d_{i}})$, yielding:
$e_{q}^{\text{FT}}$ $=C_{1}\!\left(\tfrac{m_{q}^{+}}{m_{q}}\mu_{1}+(1-\tfrac{m_{q}^{+}}{m_{q}})\mu_{2}\right)$ (9)
and similarly for SP and DP:
$e_{q}^{\text{SP}}=C_{2}(\gamma\mu_{1}+(1-\gamma)\mu_{2}),\quad e_{q}^{\text{DP}}$ $=C_{3}((\rho s(\beta\!-\!1)+m_{q}^{+})\mu_{1}$
$\quad+((1\!-\!\rho)s(\beta\!-\!1)+m_{q}\!-\!m_{q}^{+})\mu_{2})$ (10)
where $C_{1},C_{2},C_{3}$ are normalization constants. By Lemma 1:
$\gamma>\tfrac{m_{q}^{+}}{m_{q}}\Leftrightarrow s(e_{q}^{\rm SP},\mu_{1})>s(e_{q}^{\rm FT},\mu_{1})\Leftrightarrow E^{SP}>E^{FT}$ (11)
and similarly for DP. Comparing SP and DP when $\gamma=\rho$:
$E^{SP}>E^{DP}\Leftrightarrow\gamma>\tfrac{m_{q}^{+}}{m_{q}}$ (12)
Theorem 1 provides a general justification for quality-based pruning: any scoring function that identifies true positives at a rate higher than the dataset’s base rate ($m_{q}^{+}/m_{q}$) will improve learned query representations. Under equal selection quality, SP outperforms DP because it completely excludes noisy samples rather than merely down-weighting them. However, as training progresses, DP can surpass SP through improved sampling quality ($\rho$) and strength ($\beta$) enabled by dynamic threshold scheduling. This motivates a two-stage approach: applying SP first to discard strong noise, followed by DP to refine on the filtered data. The empirical validation is presented in Section 3.5.
3 Experiments
To evaluate our proposed OPERA approach, we design a series of experiments addressing six key research questions:
  • RQ1: How does OPERA compare to FT and other pruning methods across domains?
  • RQ2: Can OPERA’s findings scale to LLM-based dense retrievers?
  • RQ3: How effective is OPERA in handling noisy training data?
  • RQ4: How does OPERA affect convergence speed and training efficiency?
  • RQ5: What is the computational overhead of OPERA, and how can it be optimized?
  • RQ6: How does DP’s sampling behavior evolve during training?
We evaluate our methods on eight datasets spanning six domains: NFCorpus (Boteva et al., 2016) (nutrition), TripClick head/torso (Rekabsaz et al., 2021) (medical), FiQA (Maia et al., 2018) (finance), ANTIQUE (Hashemi et al., 2020) (non-factoid QA), TriviaQA (Joshi et al., 2017) and HotpotQA (Yang et al., 2018) (factoid QA), and FEVER (Thorne et al., 2018) (fact verification). FEVER and HotpotQA were seen during bge-large-en-v1.5 pretraining (Xiao et al., 2023), while the others are unseen. See Appendix C.1 for more datasets statistics.
3.1 Implementation Details
We used the bge-large-en-v1.5 model (Xiao et al., 2023) (335M parameters) as our primary dense retriever. All hyperparameters were selected to optimize the FT baseline performance and used without any additional tuning for all other methods, ensuring a fair comparison. We also compare against Random Pruning (RP), which retains the same fraction of data as SP but selects pairs randomly; RP consistently underperforms all other methods and is therefore excluded from the main results table but included in the efficiency analysis (Figure 1) and the full results in Appendix E. Dataset-specific hyperparameters are detailed in Appendix C.2. For LLM-based retriever experiments (Section 3.2.2), we use Qwen3-Embedding-0.6B with hyperparameters similarly optimized for its FT baseline; implementation details are provided in Appendix C.3.
3.2 Main Results
Table 1: OPERA vs. baselines on bge-large-en-v1.5. Best in bold, second-best underlined. Datasets seen during pretraining.
NDCG@10 Recall@20
OPERA OPERA
Domain Dataset PT FT IB SP DP PT FT IB SP DP
Nutrition NFCorpus 0.451 0.466 0.470 0.491 0.480 0.232 0.300 0.304 0.267 0.304
Medical TripClick (h) 0.219 0.298 0.295 0.270 0.309 0.189 0.242 0.244 0.219 0.252
TripClick (t) 0.208 0.245 0.248 0.236 0.249 0.334 0.398 0.396 0.373 0.400
Finance FiQA 0.489 0.514 0.514 0.516 0.524 0.602 0.639 0.640 0.630 0.639
Non-Factoid QA ANTIQUE 0.543 0.575 0.572 0.564 0.590 0.414 0.405 0.398 0.428 0.413
Factoid QA TriviaQA 0.487 0.481 0.482 0.501 0.491 0.429 0.458 0.460 0.445 0.460
HotpotQA 0.790 0.806 0.806 0.803 0.812 0.807 0.836 0.835 0.781 0.838
Fact Verif. FEVER 0.868 0.892 0.893 0.915 0.902 0.950 0.960 0.961 0.950 0.962
Average 0.507 0.535 0.535 0.537 0.545 0.495 0.530 0.530 0.512 0.534
Avg. Rank (Unseen) 4.67 3.33 3.00 2.67 1.33 4.50 2.83 2.33 3.50 1.83
Avg. Rank (Seen) 5.00 3.50 2.50 2.50 1.50 4.00 2.50 2.50 5.00 1.00
Avg. Rank (Overall) 4.75 3.38 2.88 2.63 1.38 4.38 2.75 2.38 3.88 1.63
PT: Pretrained, IB: InfoBatch, SP: Static Pruning, DP: Dynamic Pruning. Datasets seen in pretraining.
Table 2: OPERA vs. baselines on Qwen3-Embedding-0.6B. Best in bold, second-best underlined.
NDCG@10 Recall@20
OPERA OPERA
Domain Dataset PT FT IB SP DP PT FT IB SP DP
Nutrition NFCorpus 0.441 0.479 0.478 0.487 0.479 0.211 0.314 0.308 0.265 0.311
Non-Factoid QA ANTIQUE 0.518 0.496 0.506 0.540 0.520 0.398 0.340 0.329 0.396 0.353
Factoid QA TriviaQA 0.467 0.489 0.484 0.493 0.504 0.403 0.462 0.453 0.421 0.461
Average 0.475 0.488 0.489 0.507 0.501 0.337 0.372 0.363 0.361 0.375
3.2.1 bge-large-en-v1.5
We conduct a comparative analysis of SP and DP against the pretrained model (Xiao et al., 2023), standard finetuning (FT) (Xiao et al., 2023), and InfoBatch (Qin et al., 2023). The evaluation metrics are NDCG (Järvelin and Kekäläinen, 2002; Thakur et al., 2021; Muennighoff et al., 2022) and Recall (Zhan et al., 2021; Chen et al., 2024), assessed at the top 10 and 20 retrievals, respectively, with all methods trained for an equivalent number of iterations.
SP confirms the quality-coverage tradeoff in the previous section: it outperforms FT and InfoBatch in NDCG@10 across unseen, seen, and all datasets on average (average rank: 2.63), achieving the best NDCG@10 on NFCorpus (0.491), TriviaQA (0.501), and FEVER (0.915). As expected from our analysis, this comes at the cost of recall, as SP’s Recall@20 average rank drops to 3.88, because a priori removal of training pairs reduces query diversity. Despite this, SP offers substantial data efficiency: it drops 75% of pairs yet improves ranking, and subsequent analysis demonstrates even faster convergence and effective denoising capabilities.
DP resolves the quality-coverage tradeoff, achieving the best performance on both metrics. It achieves the highest average rank on NDCG@10 (1.38) and Recall@20 (1.63), consistently outperforming other methods for both unseen (NDCG@10: 1.33, Recall@20: 1.83) and seen (NDCG@10: 1.50, Recall@20: 1.00) datasets. DP achieves the highest NDCG@10 on 6 of 8 datasets and the highest Recall@20 on 5 of 8. By replacing hard exclusions, DP maintains the broad data coverage needed for recall while concentrating training effort on informative instances. An ablation of DP’s hierarchical design is presented in Section 3.3, with additional analysis on the static pruning data retention rate in Appendix D.2.
3.2.2 Qwen3-Embedding-0.6B
To investigate whether OPERA generalizes beyond encoder-only models, we evaluate on Qwen3-Embedding-0.6B (Zhang et al., 2025), a decoder-based LLM embedding model that employs last-token pooling and instruction-based query encoding, representing a fundamentally different architecture from the CLS-pooling encoder model used above. Due to the higher computational cost and potential data leakage from large-scale pretraining, we use a higher learning rate (1e-5 vs. 1e-6) with only 2,000 iterations (vs. 8,000–32,000 for BGE). Note that this setting is inherently less favorable to DP, which benefits from more iterations to dynamically adjust sampling rates. We evaluate on datasets with fewer than 1M documents (NFCorpus, ANTIQUE), with the exception of TriviaQA (21M documents), included to demonstrate scalability. We exclude FiQA, as the pretrained Qwen3-Embedding model already outperforms all finetuned methods on this dataset (NDCG@10: 0.511, Recall@20: 0.633), likely due to high-quality financial domain data in its pretraining corpus. As with the BGE experiments, all hyperparameters were first optimized for the vanilla FT baseline, and OPERA’s pruning methods were applied without additional tuning.
Table 2 presents the results. Despite the limited training budget, SP achieves the best average NDCG@10 (0.507), while DP achieves the best average Recall@20 (0.375), reproducing the same quality-coverage pattern observed with bge-large-en-v1.5: SP excels at ranking due to its focus on high-quality examples, while DP maintains stronger recall through soft pruning. Notably, even under conditions unfavorable to dynamic pruning (few iterations, high learning rate), DP still outperforms baselines on average, achieving NDCG gains similar to SP while preserving and improving recall, confirming that OPERA’s benefits extend to LLM-based retrievers. Detailed results are provided in Appendix F.
3.3 Ablation: Hierarchical Query-Document Pruning
Table 3: Ablation of hierarchical pruning. Recall@20 is 0.639 for all; Recall@10 shown to differentiate.
Method NDCG@10 Recall@10
FT 0.514 0.547
DP w/ Query Sel. 0.517 0.556
DP w/ Doc Sel. 0.516 0.549
DP w/ Both 0.524 0.559
Table 4: Computational overhead of DP with varying query update intervals ($I_{u}$) on FiQA.
Method $\mathbf{I_{u}}$ Iters/sec Speed Diff (%) NDCG@10 Recall@20
FT 2.43 0.00 0.514 0.639
InfoBatch 2.42 -0.30 0.514 0.640
DP 1 2.32 -4.49 0.524 0.639
DP 10 2.37 -2.51 0.519 0.646
DP 100 2.39 -1.64 0.522 0.635
A key design choice in DP is operating at two granularities, query selection and document selection, reflecting the two-stage sampling structure of retrieval training. To validate this design, we ablate each component on FiQA (Maia et al., 2018) (Table 4). Query selection alone (NDCG@10: 0.517, Recall@10: 0.556) and document selection alone (NDCG@10: 0.516, Recall@10: 0.549) both improve over FT (NDCG@10: 0.514, Recall@10: 0.547), confirming that both granularities carry complementary signal. Their combination achieves the highest scores on both metrics (NDCG@10: 0.524, Recall@10: 0.559). This supports our core argument that retrieval-specific data pruning must account for the query-document hierarchy. An additional ablation on the static pruning data retention rate is provided in Appendix D.2.
3.4 Efficiency Analysis
3.4.1 Convergence Speed
Figure 1: Training efficiency on ANTIQUE (unseen) and FEVER (seen). RP and SP use retention rate k=0.25k{=}0.25.
Figure 1 illustrates the training efficiency on both the unseen ANTIQUE dataset (Hashemi et al., 2020) and the FEVER dataset (Thorne et al., 2018) seen during pretraining. We evaluate SP and DP against FT and alternative pruning approaches across multiple training iterations. To ensure a fair comparison, we conducted separate experiments for each iteration count without using checkpoints, maintaining full learning rate scheduling in all experiments. Note that for DP, each run produces a fundamentally different sampling trajectory due to the dependence of the cosine schedule on $t_{max}$, yet DP consistently outperforms baselines across all iteration counts, demonstrating robustness to these scheduling variations.
Both pruning strategies demonstrate substantial efficiency gains. On FEVER, while FT and InfoBatch (Qin et al., 2023) require 16,000 iterations to achieve optimal NDCG@10, DP achieves comparable results in fewer than 8,000 iterations, and SP reaches this level in fewer than 500 iterations. SP shows particularly rapid convergence in early training stages and consistently outperforms baseline approaches on NDCG@10, while DP demonstrates robustness across both NDCG@10 and Recall@20. Although DP introduces a small per-iteration overhead (4.5%), this is offset by requiring fewer than 50% of the training iterations to reach peak performance, making DP more efficient overall.
3.4.2 Computation Overhead
Table 4 compares the computational costs of DP with varying query update intervals ($I_{u}$). With the most frequent updates ($I_{u}=1$), DP processes 2.32 iterations per second compared to FT’s 2.43, a 4.49% reduction. Setting $I_{u}=100$ reduces this to 1.64% while maintaining comparable performance (NDCG@10 = 0.522, Recall@20 = 0.635), and $I_{u}=10$ achieves the highest Recall@20 (0.646) with only 2.51% overhead. The document update interval has negligible computational impact.
Table 5: Denoising evaluation on ANTIQUE with noisy positive samples. SPk: retention rate $k$; SPk+DP: two-stage pipeline. Best in bold, second-best underlined.
Metric Baselines OPERA (Individual) OPERA (Two-Stage)
Pretrained FT InfoBatch SP.25 SP.5 SP.75 DP SP.25+DP SP.5+DP SP.75+DP
NDCG@10 0.543 0.570 0.567 0.560 0.560 0.582 0.587 0.560 0.567 0.582
Recall@20 0.414 0.395 0.390 0.430 0.430 0.419 0.411 0.426 0.433 0.422
3.5 Denoising Capability
We evaluate OPERA’s robustness to label noise by introducing noisy samples into the ANTIQUE (Hashemi et al., 2020) training set. Specifically, we include documents with lower relevance levels (level 2, which “does not answer the question”) as positives, simulating real-world scenarios with imperfect annotations. The detailed setup is provided in Appendix C.4.
Table 5 presents the results. Consistent with Theorem 1, which predicts that pruning outperforms FT when selection quality exceeds random, a clear divergence emerges between ranking and retrieval metrics under noisy conditions: while NDCG@10 improves from 0.543 (Pretrained) to 0.570 (FT), Recall@20 degrades from 0.414 to 0.395, indicating that retrieval effectiveness is more sensitive to false-positive training signals than ranking performance.
Among individual methods, DP achieves the highest NDCG@10 (0.587), while SP yields substantially better Recall@20 (0.430) than both FT (0.395) and DP (0.411). This advantage of SP in noisy settings motivates the two-stage approach: applying SP first to filter out noisy samples, followed by DP on the filtered data. The combined SP.5+DP achieves the best overall Recall@20 (0.433). Notably, SP.75+DP outperforms all baselines on both metrics (NDCG@10: 0.582, Recall@20: 0.422), confirming that OPERA provides an effective denoising mechanism for dense retriever finetuning.
3.6 Sampling Weight Visualization
To analyze how DP allocates training focus over time, we visualize the evolution of sampling probabilities on the FiQA (Maia et al., 2018) dataset (Figure 2). Unlike curriculum learning (Bengio et al., 2009), which imposes a fixed easy-to-hard ordering and may exclude examples at certain stages, DP keeps all examples accessible throughout training while continuously re-evaluating and adjusting their sampling weights as model representations evolve.
Figure 2(a) shows the probability distribution grouped by query. Crucially, nearly all queries maintain nonzero sampling probabilities throughout training, demonstrating that DP preserves broad query coverage unlike SP, which would exclude many queries entirely. The variation in intensity reflects quality-aware up-weighting while maintaining the diversity needed for recall. Figure 2(b), sorted by initial probabilities, reveals how DP dynamically redistributes attention: initially high-probability examples may decrease in importance while previously low-priority examples gain prominence. This demonstrates DP’s ability to discover valuable training examples beyond what the pretrained model initially favors, shifting focus toward examples more informative for domain-specific adaptation.
(a) Grouped by query
(b) Sorted by initial prob.
Figure 2: DP sampling probability evolution. Color: black (low) to yellow (high).
4 Conclusion
We presented OPERA, a data pruning framework for domain adaptation of dense retrievers. Our investigation revealed a quality-coverage tradeoff intrinsic to the two-stage query-document sampling structure of retrieval training: static pruning (SP) improves ranking by focusing on high-quality pairs but reduces retrieval coverage, while dynamic pruning (DP) resolves this tradeoff through soft sampling modulation, achieving the best performance on both ranking and retrieval metrics across most evaluation settings while halving convergence time. Both approaches demonstrate effective denoising capability, especially when combined in a two-stage pipeline. Experiments on Qwen3-Embedding-0.6B provide evidence that these findings generalize beyond encoder-only architectures.
In practice, the choice depends on the application: SP is suited for ranking-focused scenarios where training speed is paramount, DP is preferred when both ranking and retrieval performance matter, and SP followed by DP is recommended when training data contains known label noise. We hope that the quality-coverage tradeoff identified in this work provides a useful lens for future research on efficient training for retrieval models.
Appendix A Related Work
A.1 Dense Retrieval Models
Recent advances in dense retrieval have demonstrated significant improvements over traditional sparse retrieval methods. These models leverage pretrained language models (Devlin et al., 2018; Jiang et al., 2023; Touvron et al., 2023) to generate robust embeddings in various retrieval tasks (Muennighoff et al., 2022; Thakur et al., 2021). While models like NV-Embed-v1 (Lee et al., 2024) with 7B parameters have pushed performance boundaries, a more compact model family such as BGE (Xiao et al., 2023; Chen et al., 2024) offers an attractive balance between computational efficiency and effectiveness. Our work is based on the bge-large-en-v1.5 model (Xiao et al., 2023), for its favorable balance between computational efficiency and strong performance, as well as its open source training and evaluation process, which enables reproducibility and practical deployment. This choice allows us to conduct extensive experiments on various data pruning strategies while maintaining manageable computational requirements. More recently, LLM-based embedding models such as Qwen3-Embedding (Zhang et al., 2025) have further advanced the state of the art by adapting large language models directly for embedding generation, achieving strong retrieval performance across diverse tasks, though at substantially higher computational cost.
A.2 Data Pruning in Neural Networks
Data pruning has emerged as a promising approach to improve training efficiency while maintaining or enhancing model performance (Yang et al., 2022; Raju et al., 2021; Toneva et al., 2018; He et al., 2024b; Dong et al., 2023). Recent innovations have shown that data pruning can surpass traditional power law scaling, improving efficiency with minimal performance degradation (Sorscher et al., 2022). Previous work in this area can be categorized into two main approaches:
Static Pruning
Traditional static pruning methods select a fixed subset of training data before the training process begins (Toneva et al., 2018; Killamsetty et al., 2021; Mirzasoleiman et al., 2020). While these approaches can reduce training time, they often struggle with generalization across different architectures and datasets. In the context of dense retrievers, (Wang et al., 2022) introduced the consistency-based filter for pretraining, which retains only high-quality text pairs based on their ranking against random documents. Our work extends these ideas to the domain adaptation setting, where different considerations apply due to the distinct nature of the training data.
Dynamic Pruning
More recent approaches have explored dynamic data selection during training (Raju et al., 2021; He et al., 2024a; Qin et al., 2023; Li et al., 2024; Huang et al., 2024; Xiao et al., 2024). InfoBatch (Qin et al., 2023) notably achieves training acceleration through unbiased dynamic pruning, adjusting sampling probabilities based on training loss. However, these methods primarily focus on maintaining model performance while improving convergence speed. In contrast, our approach demonstrates that careful selection of query-document pair can simultaneously enhance both training efficiency and model performance.
To the best of our knowledge, there is no prior work discussing data pruning for dense retrievers during the finetuning stage, where the query is selected regardless of its relevant documents. Therefore, we will compare our approach with our implementation of InfoBatch (Qin et al., 2023) which prunes on the query level.
A.3 Curriculum Learning
Curriculum learning has emerged as a promising approach to train neural networks by presenting training examples in a meaningful order. The idea of curriculum learning was formalized in (Bengio et al., 2009), where it showed that gradually increasing the difficulty of training examples could lead to better generalization and faster convergence. In natural language processing, CL has been applied to machine translation (Platanios et al., 2019), question answering (Sachan and Xing, 2016), language modeling (Graves et al., 2017), etc. Several approaches have been proposed for automatically determining the difficulty of training examples and generating curricula. Self-paced learning (Kumar et al., 2010) allows the model to automatically select its own curriculum based on the loss of training examples. Other works have explored the use of teacher-student frameworks (Matiisen et al., 2019), where a teacher model determines the curriculum for a student model.
Our dynamic pruning framework incorporates curriculum learning principles through its evolving threshold scheduling, which gradually adjusts both the ratio and sampling strength of training instances as the model’s representations become more refined. However, a key distinction is that OPERA selects examples based on data quality (query-document similarity) rather than difficulty. These are distinct axes: a high-similarity query-document pair can be easy or hard depending on the negative mining strategy, and a difficult example is not necessarily high-quality. This quality-based perspective, combined with the hierarchical query-document sampling structure specific to retrieval, differentiates our approach from standard curriculum learning.
Appendix B Theoretical Analysis
The full theoretical analysis, including Lemma 1, Theorem 1, and their proofs, is presented in Section 2.4 of the main paper.
Appendix C Experimental Setup
C.1 Dataset Statistics
Table 6 presents detailed statistics for all eight evaluation datasets. The datasets vary significantly in scale, ranging from NFCorpus with 3,633 documents to TriviaQA with over 21 million documents. Training set sizes span from 2,426 queries (ANTIQUE) to 109,810 queries (FEVER), while the number of positive query-document pairs ranges from 14,166 (FiQA) to 741,436 (TriviaQA), reflecting diverse annotation densities across domains. FEVER and HotpotQA were previously seen by the pretrained model during its initial training, while the remaining six datasets are unseen.
Table 6: Dataset statistics. #pos: number of positive query-document pairs. Datasets seen during pretraining.
Dataset Domain #Docs Train Test
#q #pos #q #pos
NFCorpus Nutrition 3,633 2,590 110,575 323 12,334
TripClick (h) Medical 1,523,878 3,529 55,663 1,175 32,067
TripClick (t) Medical 1,523,878 105,964 424,820 1,175 6,202
FiQA Finance 57,638 5,500 14,166 648 1,706
ANTIQUE Non-Factoid QA 403,666 2,426 19,813 200 2,976
TriviaQA Factoid QA 21,015,324 78,785 741,436 8,837 82,658
HotpotQA Factoid QA 5,233,329 85,000 170,000 7,405 14,810
FEVER Fact Verif. 5,416,568 109,810 140,085 6,666 7,937
Datasets seen by the pretrained model.
C.2 Dataset-specific Hyperparameters
Table 7 summarizes the dataset-specific hyperparameters used during training. All hyperparameters were optimized based on the vanilla FT baseline performance.
Table 7: Dataset-specific hyperparameters, optimized for the FT baseline.
Dataset Negative Mining $\mathbf{k}$ Training
Mining Range Iters
NFCorpus Random 0.25 32,000
TripClick (h) Hard 100–1,100 0.25 32,000
TripClick (t) Hard 100–1,100 0.25 32,000
FiQA Hard 10–100 0.50 8,000
TriviaQA Random 0.25 16,000
ANTIQUE Hard 50–500 0.25 16,000
FEVER Hard 10–200 0.25 32,000
HotpotQA Hard 10–200 0.25 32,000
The model finetuning process uses a learning rate of 1e-6 with a linear scheduler and operates on a per-device train batch size of 8, yielding a total batch size of 64. Training uses FP16 precision with a maximum gradient norm of 1.0, no warmup, and no weight decay. The temperature is set to 0.02, and the system leverages cross-device negatives during training. The maximum query and passage length are set to 128 and 512 tokens, respectively. All embeddings are normalized. FiQA (Maia et al., 2018)’s data retention ratio $k$ is set to 0.5 to accommodate its smaller dataset size.
For dynamic pruning, we initialize the query cutoff ratio ($r_{s}$) at 0.25, and use sampling strengths of 2 and 5 for the starting ($\alpha_{s}$) and ending ($\alpha_{e}$) values, respectively. The document level includes an initial cutoff ratio ($v_{s}$) of 0.25, a terminal cutoff ratio ($v_{e}$) of 0.5, and a constant sampling strength ($\beta$) of 5. All cutoff schedulers follow a cosine schedule.
C.3 Qwen3-Embedding-0.6B Implementation Details
[…]
C.4 Denoising Experiment Setup
The ANTIQUE dataset (Hashemi et al., 2020) defines relevance levels as follows:
  • Level 1: Completely out of context or does not make any sense (4.6% of the training data)
  • Level 2: Does not answer the question, or provides an unreasonable answer, but is not out of context (23.1% of the training data)
  • Level 3: Can be an answer to the question, but is not sufficiently convincing (29.5% of the training data)
  • Level 4: Looks reasonable and convincing, with high quality (42.8% of the training data)
Our experimental setup is designed as follows:
  • Test set: We maintain the standard evaluation criterion, considering only documents with relevance levels of 3 and 4 as positive samples (Hashemi et al., 2020).
  • Training set: Documents with relevance levels of 2, 3, and 4 were treated as positive samples. The inclusion of level 2 documents, which are documented as insufficient answers, deliberately introduces noise into the positive samples. We did not include level 1 documents as positive samples since they represent a small portion (4.6%) of the training data and are more like random noise, which differs from real-world scenarios where noisy samples are typically hard negatives that share some relevance with the query.
Appendix D Additional Experiments
D.1 Consistency-Based Score Analysis
Table 8: SP with cosine similarity vs. CBS as pruning metric. Best per row in bold.
Dataset Metric SP SP (CBS)
NFCorpus NDCG@10 0.491 0.463
Recall@20 0.267 0.305
FiQA NDCG@10 0.511 0.514
Recall@20 0.620 0.635
ANTIQUE NDCG@10 0.564 0.577
Recall@20 0.428 0.403
FEVER NDCG@10 0.915 0.894
Recall@20 0.950 0.961
Average NDCG@10 0.620 0.612
Recall@20 0.566 0.576
The consistency-based filter was introduced in (Wang et al., 2022) as a quality control mechanism for large-scale pretraining. The core insight is that high-quality training pairs should maintain relevance compared to random documents: the method ranks each positive pair against random negatives and retains only those that consistently rank highly. We adapt this to our finetuning context with three modifications: (1) we leverage the pretrained checkpoint directly instead of training on noisy data, (2) we reduce the random document pool from one million to ten thousand to match the smaller scale of downstream tasks, and (3) we replace the static top-k threshold with a reciprocal rank metric (CBS) for percentage-based filtering. We sample random negatives from positive documents of other queries to avoid additional embedding computation.
We compare SP using CBS with SP using cosine similarity. Cosine similarity directly measures the semantic relationship between queries and documents, while CBS evaluates pairs based on their relative ranking against random samples.
Table 8 presents the comparative performance across these datasets. SP uses cosine similarity as the similarity metric for pruning, while SP (CBS) uses CBS. In the nutrition dataset NFCorpus, SP achieves optimal NDCG@10 (0.491), while SP (CBS) leads in Recall@20 (0.305). FiQA demonstrates marginal improvements with SP (CBS) on both metrics (NDCG@10: 0.514, Recall@20: 0.635). For ANTIQUE, SP (CBS) leads in NDCG@10 (0.577), while SP with cosine similarity achieves higher Recall@20 (0.428). In FEVER, SP excels in NDCG@10 (0.915), while SP (CBS) shows superior Recall@20 (0.961). Overall, SP achieves better NDCG@10 (0.620) while SP (CBS) performs better on Recall@20 (0.576).
However, CBS is incompatible with DP as it requires computationally expensive recalculation of embeddings on all negative documents after each model update. We therefore adopt cosine similarity as the default pruning metric for both SP and DP.
D.2 Static Pruning Data Retention Rate
Figure 3: Effect of SP data retention rate $k$ on NFCorpus. SP improves NDCG@10 over FT at all retention rates.
We evaluate the effects of varying data retention rates ($k$) in Static Pruning (SP) through comprehensive experiments on NFCorpus (Boteva et al., 2016). The experimental setup comprises six configurations: standard finetuning (FT) as baseline ($k=1$) and SP with $k=0.75,0.5,0.25,0.1,0.05$.
The results, presented in Figure 3, demonstrate that SP consistently surpasses FT in terms of NDCG@10 across all retention rates. The optimal performance was achieved at $k=0.25$, achieving a peak NDCG of 0.491. Notably, even with minimal data retention (5%), SP maintains superior performance with an NDCG of 0.482 compared to FT (0.466). As expected, Recall@20 shows a monotonic decrease as the retention rate reduces from 100% to 5%, which aligns with the reduced data visibility. These findings highlight SP’s data efficiency and indicate the potential for computational resource optimization while maintaining or enhancing ranking performance.
Appendix E Detailed bge-large-en-v1.5 Results
We present full evaluation results for bge-large-en-v1.5 across all eight datasets at multiple training iterations (0 to 32,000). Each table reports NDCG, Recall, and Success at cutoffs of {10, 20, 100} for all methods: FT, RP (Random Pruning with the same retention rate as SP), InfoBatch, SP, and DP. The subscript in RP and SP denotes the data retention rate $k$ (e.g., RP25 retains 25% of training pairs selected randomly, SP25 retains the top 25% by similarity). FiQA additionally includes SP50 and RP50 variants due to its smaller dataset size. Results are shown for NFCorpus (Table 9), FiQA (Table 10), ANTIQUE (Table 11), TriviaQA (Table 12), TripClick head (Table 13), TripClick torso (Table 14), FEVER (Table 15), and HotpotQA (Table 16).
Appendix F Detailed Qwen3-Embedding-0.6B Results
We present comprehensive evaluation results for Qwen3-Embedding-0.6B across four datasets at 2,000 training iterations: NFCorpus (Table 17), FiQA (Table 18), ANTIQUE (Table 19), and TriviaQA (Table 20). We report all evaluation metrics (MRR, NDCG, Recall, and Success at cutoffs of {1, 5, 10, 20, 50, 100}) for each configuration. The pretrained model results are iteration-independent and shown as a reference in each table. Notably, for FiQA (Table 18), the pretrained model outperforms all finetuned methods across all metrics, suggesting that Qwen3-Embedding’s pretraining corpus already contains high-quality financial domain data.
Appendix G Limitations
We identify several directions for future work. Our LLM-based experiments use Qwen3-Embedding-0.6B as initial evidence; extending to larger models (4B, 8B parameters) and additional model families would strengthen the generalizability claim, though we note that the consistent quality-coverage tradeoff pattern across two architecturally distinct models (encoder-only BGE and decoder-based Qwen3) suggests the underlying principle is architecture-agnostic. Our evaluation covers eight English-language text retrieval datasets across six domains; applying OPERA to multilingual and multimodal retrieval settings is a natural extension. While we report results under identical hyperparameters (optimized for the FT baseline of BGE) but use a higher learning rate for Qwen3 due to the substantially higher computational cost of LLM-based retrievers, we do not report variance across random seeds due to the computational cost of running all method-dataset combinations multiple times; however, the consistency of improvements across eight diverse datasets provides evidence of robustness beyond what single-dataset variance estimates would capture. Finally, OPERA’s dynamic pruning hyperparameters (sampling strengths, cutoff ratios, and scheduling) were not tuned due to computational constraints; exploring different configurations or adaptive scheduling strategies may yield further gains.
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.45101 0.45101 0.45101 0.45101 0.45101
500 0.45608 0.45768 0.46132 0.47209 0.46939
1000 0.46758 0.46583 0.46669 0.47429 0.47479
2000 0.47059 0.46631 0.47209 0.47979 0.47418
4000 0.47465 0.47516 0.47438 0.48804 0.48388
8000 0.47471 0.45486 0.47866 0.49103 0.49008
16000 0.4802 0.43449 0.48371 0.48671 0.48067
32000 0.46565 0.42648 0.46965 0.49128 0.48019
Recall@10 0 0.18786 0.18786 0.18786 0.18786 0.18786
500 0.19205 0.19057 0.1934 0.19968 0.19693
1000 0.20346 0.19951 0.20231 0.198 0.20538
2000 0.20952 0.20231 0.20912 0.20422 0.20623
4000 0.21562 0.20594 0.21102 0.20835 0.21765
8000 0.21083 0.20026 0.21204 0.21186 0.21774
16000 0.21814 0.18932 0.22421 0.20753 0.22189
32000 0.22654 0.18213 0.23186 0.21049 0.2272
Success@10 0 0.75542 0.75542 0.75542 0.75542 0.75542
500 0.74923 0.73994 0.74613 0.75851 0.75851
1000 0.75542 0.74613 0.75232 0.75542 0.76471
2000 0.76161 0.75542 0.76161 0.76471 0.76161
4000 0.7709 0.77709 0.76471 0.77709 0.77709
8000 0.77709 0.76471 0.78019 0.78019 0.7678
16000 0.76471 0.73684 0.76161 0.76471 0.78019
32000 0.75851 0.72446 0.74923 0.75232 0.74613
NDCG@20 0 0.46963 0.46963 0.46963 0.46963 0.46963
500 0.47225 0.47045 0.47562 0.48777 0.4863
1000 0.48135 0.47612 0.48132 0.493 0.49069
2000 0.48129 0.48145 0.4836 0.50005 0.48827
4000 0.48389 0.48097 0.48472 0.50418 0.49289
8000 0.48058 0.46252 0.48539 0.50196 0.50145
16000 0.48987 0.44691 0.49305 0.50082 0.49279
32000 0.476 0.43441 0.48058 0.4989 0.49261
Recall@20 0 0.23235 0.23235 0.23235 0.23235 0.23235
500 0.24643 0.23986 0.24754 0.24674 0.25311
1000 0.25214 0.24472 0.2516 0.25396 0.25597
2000 0.2572 0.25875 0.25863 0.25857 0.25766
4000 0.2667 0.25901 0.26489 0.26086 0.26595
8000 0.27199 0.25249 0.27128 0.25804 0.27595
16000 0.29227 0.24643 0.29663 0.26297 0.292
32000 0.30031 0.24182 0.30361 0.2671 0.30398
Success@20 0 0.80495 0.80495 0.80495 0.80495 0.80495
500 0.79876 0.79567 0.80186 0.80186 0.81115
1000 0.79567 0.78947 0.79567 0.81734 0.80805
2000 0.79876 0.80495 0.79876 0.81734 0.80186
4000 0.81734 0.80805 0.80805 0.80805 0.81424
8000 0.81115 0.80805 0.82353 0.81424 0.81424
16000 0.82353 0.78328 0.82663 0.81115 0.81734
32000 0.79567 0.76161 0.79257 0.80186 0.80805
NDCG@100 0 0.5908 0.5908 0.5908 0.5908 0.5908
500 0.59424 0.59384 0.59712 0.60724 0.60158
1000 0.60187 0.5992 0.60232 0.60983 0.60755
2000 0.60681 0.60372 0.60681 0.61762 0.6134
4000 0.61242 0.61026 0.61227 0.62192 0.62047
8000 0.61675 0.5967 0.62023 0.6202 0.62778
16000 0.61673 0.58437 0.61839 0.62012 0.61766
32000 0.60532 0.57457 0.60861 0.61819 0.61702
Recall@100 0 0.36386 0.36386 0.36386 0.36386 0.36386
500 0.39106 0.39251 0.39036 0.38694 0.39014
1000 0.40479 0.4013 0.40445 0.3935 0.40153
2000 0.4134 0.40905 0.41078 0.39942 0.41369
4000 0.43184 0.42072 0.43176 0.40591 0.4305
8000 0.46136 0.43142 0.46095 0.40966 0.45188
16000 0.48035 0.43325 0.47794 0.41957 0.47781
32000 0.50295 0.42943 0.49731 0.42292 0.49686
Success@100 0 0.89474 0.89474 0.89474 0.89474 0.89474
500 0.89783 0.89474 0.89783 0.89474 0.88854
1000 0.89164 0.89474 0.89474 0.89783 0.88854
2000 0.90093 0.89783 0.89474 0.90402 0.89783
4000 0.90402 0.90093 0.89474 0.89783 0.90712
8000 0.89474 0.89474 0.89474 0.88545 0.89783
16000 0.88235 0.89164 0.87926 0.89783 0.89783
32000 0.88854 0.88854 0.88545 0.89164 0.89783
Table 9: Full Results on NFCorpus.
Metric #iterations FT RP50 InfoBatch SP25 SP50 DP
NDCG@10 0 0.48919 0.48919 0.48919 0.48919 0.48919 0.48919
500 0.5073 0.50502 0.50971 0.50296 0.50469 0.50404
1000 0.5085 0.50659 0.50761 0.50241 0.50876 0.50655
2000 0.51277 0.50653 0.51195 0.50229 0.51293 0.51212
4000 0.5148 0.5026 0.51559 0.51047 0.51555 0.51812
8000 0.51359 0.50572 0.5137 0.511 0.5162 0.52418
16000 0.50859 0.49985 0.50758 0.51213 0.51463 0.51715
Recall@10 0 0.51432 0.51432 0.51432 0.51432 0.51432 0.51432
500 0.53982 0.53969 0.53958 0.53143 0.5367 0.54259
1000 0.54812 0.5451 0.54982 0.53329 0.54211 0.54512
2000 0.559 0.54228 0.55506 0.53052 0.54561 0.5546
4000 0.55662 0.54005 0.55558 0.54067 0.54815 0.56327
8000 0.54711 0.53811 0.54964 0.53731 0.54906 0.55887
16000 0.54124 0.52443 0.53852 0.54046 0.5469 0.55478
Success@10 0 0.72685 0.72685 0.72685 0.72685 0.72685 0.72685
500 0.74383 0.74383 0.74691 0.73611 0.74228 0.74846
1000 0.75154 0.74537 0.75 0.74074 0.74691 0.74846
2000 0.7608 0.74383 0.75772 0.73457 0.74846 0.7608
4000 0.76235 0.75 0.76389 0.74074 0.75 0.76543
8000 0.75463 0.75154 0.75926 0.73457 0.74537 0.76389
16000 0.74691 0.74074 0.74537 0.73765 0.74383 0.76389
NDCG@20 0 0.52218 0.52218 0.52218 0.52218 0.52218 0.52218
500 0.54164 0.5384 0.54391 0.53541 0.53779 0.53734
1000 0.54275 0.54106 0.54071 0.53541 0.54142 0.54203
2000 0.54417 0.54016 0.54539 0.53641 0.54587 0.54416
4000 0.54578 0.53642 0.54629 0.54136 0.54946 0.54855
8000 0.54723 0.54071 0.54671 0.54296 0.54924 0.5539
16000 0.54209 0.53193 0.54116 0.54289 0.54821 0.54795
Recall@20 0 0.60161 0.60161 0.60161 0.60161 0.60161 0.60161
500 0.63196 0.62872 0.63162 0.61578 0.62241 0.63024
1000 0.6369 0.63292 0.6364 0.61609 0.62523 0.63847
2000 0.63903 0.62874 0.64027 0.61811 0.62882 0.63727
4000 0.63808 0.63194 0.63514 0.62047 0.63412 0.64222
8000 0.63933 0.63279 0.63994 0.61994 0.63041 0.63909
16000 0.62971 0.61118 0.63012 0.61941 0.63243 0.63809
Success@20 0 0.79321 0.79321 0.79321 0.79321 0.79321 0.79321
500 0.81636 0.81327 0.81636 0.80093 0.81019 0.81481
1000 0.82407 0.82253 0.82407 0.80401 0.81019 0.82716
2000 0.82562 0.81327 0.82407 0.80556 0.81944 0.82562
4000 0.82099 0.81481 0.8179 0.8071 0.82407 0.82407
8000 0.8287 0.81327 0.82562 0.80864 0.81944 0.82099
16000 0.8179 0.80247 0.81481 0.8071 0.8179 0.82253
NDCG@100 0 0.57052 0.57052 0.57052 0.57052 0.57052 0.57052
500 0.58557 0.58211 0.58748 0.57741 0.58011 0.58142
1000 0.58628 0.58487 0.58404 0.57796 0.58414 0.58462
2000 0.58806 0.58638 0.58894 0.57922 0.58792 0.58797
4000 0.59107 0.58097 0.59279 0.58342 0.59066 0.59249
8000 0.59063 0.58115 0.5907 0.58538 0.59152 0.59951
16000 0.58399 0.57598 0.58324 0.58426 0.58982 0.59
Recall@100 0 0.77048 0.77048 0.77048 0.77048 0.77048 0.77048
500 0.78785 0.78461 0.78634 0.75805 0.76858 0.78739
1000 0.79259 0.79034 0.79133 0.76416 0.77228 0.7897
2000 0.7951 0.7905 0.79659 0.76461 0.77901 0.79448
4000 0.79596 0.78875 0.79938 0.76684 0.78161 0.79421
8000 0.79368 0.77385 0.79628 0.76688 0.78176 0.80139
16000 0.7764 0.76227 0.77643 0.76078 0.77788 0.78376
Success@100 0 0.89506 0.89506 0.89506 0.89506 0.89506 0.89506
500 0.90586 0.89969 0.90432 0.89352 0.89815 0.90586
1000 0.90895 0.90586 0.90586 0.89506 0.89969 0.90432
2000 0.90895 0.90741 0.90895 0.89352 0.90432 0.90586
4000 0.90895 0.90278 0.91204 0.89815 0.90432 0.90741
8000 0.91049 0.8966 0.91358 0.89969 0.90432 0.91358
16000 0.90123 0.8966 0.89969 0.89506 0.90278 0.90586
Table 10: Full Results on FiQA.
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.54361 0.54361 0.54361 0.54361 0.54361
500 0.51341 0.50884 0.51514 0.555 0.54143
1000 0.53297 0.52488 0.53106 0.56221 0.54905
2000 0.55202 0.54677 0.55086 0.56385 0.56028
4000 0.57041 0.56046 0.5673 0.56475 0.5814
8000 0.5756 0.55501 0.56865 0.56304 0.58932
16000 0.57495 0.55635 0.57182 0.56404 0.58987
32000 0.5704 0.54994 0.56128 0.56153 0.5803
Recall@10 0 0.3187 0.3187 0.3187 0.3187 0.3187
500 0.30215 0.30074 0.30394 0.32731 0.3211
1000 0.31274 0.30605 0.31297 0.33135 0.32435
2000 0.31647 0.3128 0.31478 0.33139 0.32999
4000 0.32797 0.31715 0.32603 0.33024 0.33619
8000 0.32685 0.30934 0.32143 0.32961 0.33503
16000 0.32654 0.30758 0.32119 0.33087 0.33564
32000 0.3201 0.303 0.31485 0.32149 0.3269
Success@10 0 0.915 0.915 0.915 0.915 0.915
500 0.915 0.915 0.91 0.925 0.92
1000 0.92 0.93 0.925 0.925 0.92
2000 0.93 0.93 0.93 0.925 0.925
4000 0.93 0.93 0.93 0.925 0.94
8000 0.94 0.925 0.935 0.925 0.935
16000 0.94 0.92 0.93 0.925 0.94
32000 0.93 0.92 0.935 0.92 0.935
NDCG@20 0 0.58938 0.58938 0.58938 0.58938 0.58938
500 0.55908 0.55891 0.56392 0.60371 0.58903
1000 0.58022 0.57683 0.57841 0.60726 0.59685
2000 0.60082 0.59966 0.59985 0.60895 0.60292
4000 0.61127 0.60587 0.61113 0.61126 0.62214
8000 0.61673 0.6027 0.61021 0.60916 0.63131
16000 0.61622 0.60262 0.61349 0.60949 0.62942
32000 0.61282 0.59419 0.60292 0.60826 0.6267
Recall@20 0 0.41377 0.41377 0.41377 0.41377 0.41377
500 0.38463 0.38752 0.38937 0.43058 0.41207
1000 0.3951 0.39427 0.39699 0.43161 0.41464
2000 0.40321 0.40246 0.4028 0.43278 0.41329
4000 0.40542 0.40033 0.40752 0.43143 0.4189
8000 0.40418 0.39583 0.39742 0.42876 0.41904
16000 0.40475 0.38942 0.3982 0.42807 0.41345
32000 0.39887 0.38136 0.38788 0.41849 0.41037
Success@20 0 0.94 0.94 0.94 0.94 0.94
500 0.935 0.935 0.93 0.945 0.93
1000 0.93 0.93 0.935 0.95 0.935
2000 0.94 0.935 0.945 0.95 0.945
4000 0.95 0.95 0.95 0.955 0.95
8000 0.95 0.95 0.955 0.95 0.955
16000 0.945 0.94 0.955 0.95 0.955
32000 0.945 0.935 0.95 0.94 0.955
NDCG@100 0 0.70963 0.70963 0.70963 0.70963 0.70963
500 0.68968 0.68743 0.69259 0.72254 0.71339
1000 0.70339 0.69928 0.70178 0.72514 0.72015
2000 0.71801 0.71431 0.71782 0.72624 0.72481
4000 0.72655 0.71896 0.72348 0.72845 0.73586
8000 0.72726 0.71366 0.7236 0.72757 0.73949
16000 0.72375 0.71325 0.72022 0.72788 0.73581
32000 0.71905 0.70656 0.70943 0.72737 0.73014
Recall@100 0 0.61225 0.61225 0.61225 0.61225 0.61225
500 0.59569 0.59606 0.60046 0.62704 0.61638
1000 0.59057 0.59225 0.59289 0.62929 0.61486
2000 0.59087 0.58942 0.58999 0.63046 0.61596
4000 0.58692 0.58056 0.58274 0.63033 0.60371
8000 0.57344 0.56642 0.5742 0.63086 0.59186
16000 0.56939 0.55861 0.5599 0.62933 0.58018
32000 0.56019 0.55254 0.55349 0.61963 0.57318
Success@100 0 0.97 0.97 0.97 0.97 0.97
500 0.97 0.975 0.965 0.975 0.98
1000 0.975 0.975 0.975 0.975 0.98
2000 0.975 0.97 0.975 0.975 0.975
4000 0.975 0.97 0.975 0.975 0.975
8000 0.975 0.97 0.975 0.975 0.975
16000 0.975 0.965 0.97 0.975 0.975
32000 0.97 0.965 0.96 0.97 0.97
Table 11: Full Results on ANTIQUE.
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.48724 0.48724 0.48724 0.48724 0.48724
500 0.48569 0.48454 0.48815 0.49358 0.49215
1000 0.48447 0.48167 0.48317 0.49497 0.49271
2000 0.48155 0.47685 0.48055 0.4958 0.49108
4000 0.47759 0.4755 0.47892 0.49768 0.49108
8000 0.47752 0.47493 0.47722 0.49818 0.48562
16000 0.48116 0.47636 0.48179 0.50124 0.49078
32000 0.48315 0.47358 0.48287 0.50933 0.49009
Recall@10 0 0.3418 0.3418 0.3418 0.3418 0.3418
500 0.35714 0.3568 0.35926 0.35052 0.35995
1000 0.35744 0.35596 0.35722 0.35049 0.3616
2000 0.35736 0.35347 0.35697 0.3521 0.36049
4000 0.35613 0.35363 0.35553 0.3541 0.36056
8000 0.3566 0.35452 0.35697 0.35376 0.35896
16000 0.3609 0.35589 0.36158 0.35759 0.36408
32000 0.36281 0.35282 0.36419 0.36243 0.36303
Success@10 0 0.81775 0.81775 0.81775 0.81775 0.81775
500 0.82485 0.82515 0.82456 0.81997 0.82751
1000 0.82219 0.82145 0.82115 0.81908 0.82751
2000 0.81893 0.81716 0.81997 0.81953 0.82485
4000 0.81834 0.81509 0.81538 0.8213 0.82352
8000 0.8176 0.81538 0.81864 0.81967 0.8213
16000 0.8216 0.81524 0.82249 0.82086 0.82263
32000 0.82234 0.81065 0.82293 0.82396 0.8213
NDCG@20 0 0.51949 0.51949 0.51949 0.51949 0.51949
500 0.51899 0.51739 0.52001 0.52728 0.52425
1000 0.51736 0.51491 0.51661 0.52742 0.52532
2000 0.51454 0.51056 0.51409 0.52841 0.52395
4000 0.51111 0.50845 0.51221 0.53023 0.52333
8000 0.5104 0.50777 0.51101 0.53049 0.51719
16000 0.51354 0.50923 0.51423 0.53244 0.52286
32000 0.51654 0.50718 0.51534 0.5389 0.52183
Recall@20 0 0.42887 0.42887 0.42887 0.42887 0.42887
500 0.45515 0.4531 0.45329 0.44183 0.4552
1000 0.45496 0.45392 0.45569 0.43807 0.45786
2000 0.45523 0.45265 0.45518 0.44009 0.45636
4000 0.45337 0.45062 0.45393 0.44354 0.45623
8000 0.45316 0.45215 0.45542 0.44278 0.4527
16000 0.45846 0.45417 0.46028 0.44532 0.4598
32000 0.46303 0.45253 0.46247 0.44889 0.4607
Success@20 0 0.85932 0.85932 0.85932 0.85932 0.85932
500 0.8645 0.86302 0.86479 0.86464 0.86686
1000 0.86361 0.86302 0.86479 0.8608 0.86672
2000 0.86154 0.86036 0.86346 0.86154 0.86627
4000 0.8605 0.85784 0.86154 0.86183 0.86405
8000 0.85962 0.85873 0.8605 0.86139 0.86065
16000 0.86124 0.85976 0.86228 0.8605 0.86405
32000 0.86686 0.85888 0.86391 0.86109 0.86391
NDCG@100 0 0.62251 0.62251 0.62251 0.62251 0.62251
500 0.62607 0.62481 0.62697 0.62901 0.62989
1000 0.62446 0.62222 0.62391 0.62902 0.63085
2000 0.62246 0.61865 0.62162 0.62995 0.62963
4000 0.61939 0.61735 0.62012 0.63217 0.62946
8000 0.62001 0.61741 0.62006 0.63178 0.62407
16000 0.62246 0.61867 0.62344 0.63372 0.62911
32000 0.62528 0.61683 0.62496 0.63913 0.62899
Recall@100 0 0.61507 0.61507 0.61507 0.61507 0.61507
500 0.65763 0.65614 0.65534 0.62793 0.65309
1000 0.65757 0.65696 0.65847 0.62328 0.6562
2000 0.65941 0.6572 0.6587 0.62496 0.65451
4000 0.65889 0.65706 0.65834 0.63125 0.65589
8000 0.66186 0.66088 0.66144 0.62854 0.65358
16000 0.66628 0.66374 0.66767 0.63196 0.66051
32000 0.67237 0.66268 0.6734 0.63364 0.66503
Success@100 0 0.91716 0.91716 0.91716 0.91716 0.91716
500 0.92604 0.92559 0.92352 0.91967 0.92382
1000 0.9247 0.92367 0.92411 0.9176 0.92485
2000 0.92441 0.92278 0.92411 0.9182 0.92337
4000 0.92367 0.92293 0.92293 0.92086 0.925
8000 0.92574 0.92337 0.92456 0.91982 0.92367
16000 0.92589 0.92322 0.92515 0.92012 0.9253
32000 0.92885 0.92204 0.92618 0.92071 0.92633
Table 12: Full Results on TriviaQA.
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.21935 0.21935 0.21935 0.21935 0.21935
500 0.24666 0.24734 0.24895 0.24686 0.25367
1000 0.25109 0.24906 0.25135 0.25058 0.25591
2000 0.25992 0.26174 0.25768 0.25393 0.26853
4000 0.28174 0.27356 0.27998 0.25811 0.28696
8000 0.29324 0.27947 0.29348 0.26045 0.30409
16000 0.30013 0.27998 0.30069 0.26549 0.30805
32000 0.29779 0.27999 0.29524 0.26966 0.30926
Recall@10 0 0.11412 0.11412 0.11412 0.11412 0.11412
500 0.13307 0.13346 0.13356 0.13029 0.13579
1000 0.13501 0.13431 0.13513 0.13227 0.13827
2000 0.13819 0.13903 0.13794 0.1331 0.14297
4000 0.14543 0.14158 0.14237 0.13445 0.14976
8000 0.15074 0.14488 0.1542 0.13591 0.15744
16000 0.15776 0.1449 0.15936 0.1378 0.1625
32000 0.15814 0.14411 0.15668 0.13863 0.16185
Success@10 0 0.63489 0.63489 0.63489 0.63489 0.63489
500 0.70638 0.70383 0.70553 0.68681 0.71319
1000 0.71489 0.71064 0.72255 0.69787 0.71915
2000 0.73277 0.73702 0.73447 0.70043 0.74128
4000 0.76255 0.75319 0.76255 0.71149 0.76766
8000 0.78638 0.75915 0.79404 0.71489 0.79064
16000 0.7966 0.75745 0.7966 0.72936 0.80085
32000 0.78894 0.7583 0.78979 0.73617 0.8
NDCG@20 0 0.26348 0.26348 0.26348 0.26348 0.26348
500 0.29366 0.29385 0.29566 0.29126 0.29826
1000 0.29637 0.29634 0.29706 0.29559 0.3002
2000 0.30455 0.30416 0.3036 0.30001 0.31348
4000 0.326 0.31997 0.32444 0.30256 0.33175
8000 0.33759 0.32329 0.33554 0.30537 0.34681
16000 0.33797 0.32148 0.33982 0.31071 0.35008
32000 0.33739 0.32248 0.33678 0.31654 0.35104
Recall@20 0 0.18949 0.18949 0.18949 0.18949 0.18949
500 0.21663 0.21695 0.21657 0.20815 0.21731
1000 0.21899 0.21913 0.22018 0.21162 0.21994
2000 0.22352 0.22182 0.22289 0.21433 0.22791
4000 0.22958 0.22733 0.22965 0.21376 0.23769
8000 0.24058 0.22914 0.23801 0.21579 0.24658
16000 0.24178 0.22593 0.24426 0.21874 0.25091
32000 0.24346 0.22538 0.2442 0.22014 0.25207
Success@20 0 0.75319 0.75319 0.75319 0.75319 0.75319
500 0.82128 0.81447 0.81872 0.79234 0.81617
1000 0.81702 0.82298 0.82128 0.8 0.81532
2000 0.82553 0.82043 0.82638 0.80426 0.84085
4000 0.84426 0.85021 0.84426 0.8017 0.85702
8000 0.86468 0.85362 0.86383 0.80936 0.87149
16000 0.86553 0.8366 0.86553 0.81702 0.86979
32000 0.86298 0.84 0.85957 0.81957 0.87064
NDCG@100 0 0.4582 0.4582 0.4582 0.4582 0.4582
500 0.49334 0.49328 0.4953 0.48779 0.49772
1000 0.49687 0.49633 0.49899 0.49213 0.50135
2000 0.50644 0.5062 0.50519 0.49567 0.51531
4000 0.52799 0.51896 0.52622 0.49817 0.53062
8000 0.53567 0.52151 0.53459 0.50174 0.54524
16000 0.53825 0.51991 0.53858 0.50578 0.5488
32000 0.53734 0.51888 0.53583 0.50892 0.54931
Recall@100 0 0.44129 0.44129 0.44129 0.44129 0.44129
500 0.48578 0.48488 0.48621 0.46503 0.48353
1000 0.49008 0.48937 0.49335 0.46952 0.49181
2000 0.49738 0.49507 0.49835 0.47295 0.50321
4000 0.50844 0.50081 0.50844 0.47397 0.51132
8000 0.51516 0.50016 0.51674 0.47688 0.52164
16000 0.5223 0.49831 0.52344 0.47718 0.53231
32000 0.52231 0.4936 0.52017 0.47535 0.53109
Success@100 0 0.91319 0.91319 0.91319 0.91319 0.91319
500 0.93617 0.93532 0.93617 0.92511 0.93362
1000 0.93702 0.93957 0.94298 0.92936 0.94128
2000 0.94468 0.94979 0.94638 0.93191 0.95404
4000 0.95149 0.95064 0.94979 0.93021 0.95404
8000 0.95149 0.94809 0.94809 0.93532 0.95234
16000 0.94894 0.94298 0.94979 0.93702 0.95574
32000 0.94894 0.94213 0.94723 0.93787 0.95234
Table 13: Full Results on TripClick (head).
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.20834 0.20834 0.20834 0.20834 0.20834
500 0.21923 0.21885 0.21742 0.2224 0.22026
1000 0.21736 0.21728 0.21675 0.22483 0.22031
2000 0.21753 0.21787 0.21744 0.22763 0.22117
4000 0.22105 0.21873 0.22282 0.22932 0.22612
8000 0.23183 0.22836 0.23209 0.22973 0.23327
16000 0.24165 0.2365 0.24048 0.23194 0.24341
32000 0.24507 0.23823 0.24798 0.23579 0.24942
Recall@10 0 0.21973 0.21973 0.21973 0.21973 0.21973
500 0.23999 0.23883 0.23861 0.23966 0.24437
1000 0.24527 0.24226 0.24388 0.24426 0.24363
2000 0.2469 0.24785 0.2479 0.24509 0.2488
4000 0.24662 0.24813 0.24984 0.24903 0.25319
8000 0.25737 0.25676 0.25413 0.25152 0.25867
16000 0.26549 0.26493 0.26218 0.25323 0.27092
32000 0.27042 0.26894 0.27346 0.2574 0.27578
Success@10 0 0.52085 0.52085 0.52085 0.52085 0.52085
500 0.54468 0.54383 0.54128 0.53787 0.54979
1000 0.54809 0.54809 0.54468 0.54723 0.54383
2000 0.55149 0.55234 0.55404 0.54809 0.55149
4000 0.55234 0.56085 0.56511 0.54979 0.56426
8000 0.58553 0.58383 0.58043 0.54979 0.57277
16000 0.59574 0.58979 0.59149 0.55745 0.59489
32000 0.59404 0.59149 0.59745 0.57277 0.61021
NDCG@20 0 0.26082 0.26082 0.26082 0.26082 0.26082
500 0.27579 0.27316 0.27365 0.2762 0.27515
1000 0.27235 0.27205 0.27208 0.27726 0.27659
2000 0.27267 0.27239 0.27311 0.28023 0.27754
4000 0.27849 0.2753 0.27812 0.28306 0.28408
8000 0.28709 0.28263 0.28792 0.28343 0.29082
16000 0.29727 0.28966 0.29574 0.28594 0.29924
32000 0.30304 0.29397 0.30349 0.28986 0.30561
Recall@20 0 0.33434 0.33434 0.33434 0.33434 0.33434
500 0.36105 0.35785 0.35804 0.35711 0.36212
1000 0.36231 0.35973 0.3617 0.35593 0.36389
2000 0.36339 0.36408 0.36735 0.35766 0.3686
4000 0.37043 0.36911 0.3716 0.36253 0.37939
8000 0.38016 0.37795 0.37853 0.36541 0.38278
16000 0.38796 0.38256 0.38589 0.36692 0.39138
32000 0.39753 0.39027 0.39605 0.37266 0.39998
Success@20 0 0.6383 0.6383 0.6383 0.6383 0.6383
500 0.67915 0.66468 0.66809 0.67149 0.67574
1000 0.67404 0.66553 0.67149 0.67319 0.6834
2000 0.67745 0.6766 0.67915 0.67404 0.6834
4000 0.69021 0.68681 0.69106 0.68 0.69872
8000 0.69447 0.69702 0.69277 0.68851 0.70043
16000 0.70468 0.70043 0.70128 0.69362 0.71064
32000 0.7183 0.71319 0.71319 0.69191 0.72
NDCG@100 0 0.36876 0.36876 0.36876 0.36876 0.36876
500 0.38618 0.38505 0.38477 0.3857 0.38506
1000 0.38433 0.38485 0.38481 0.38712 0.38806
2000 0.38651 0.38577 0.38586 0.39046 0.3905
4000 0.39201 0.38943 0.39245 0.39179 0.39496
8000 0.3995 0.39603 0.40126 0.39208 0.40289
16000 0.41015 0.40399 0.41008 0.39419 0.4117
32000 0.41662 0.40836 0.41801 0.39733 0.41867
Recall@100 0 0.60293 0.60293 0.60293 0.60293 0.60293
500 0.64261 0.64128 0.64145 0.63196 0.64077
1000 0.64984 0.64984 0.64989 0.63542 0.64843
2000 0.66019 0.65785 0.65856 0.6394 0.66177
4000 0.66304 0.66648 0.66787 0.6402 0.66628
8000 0.67083 0.67221 0.6732 0.64336 0.67385
16000 0.68717 0.68287 0.6885 0.64703 0.68699
32000 0.69987 0.68954 0.70005 0.64903 0.69874
Success@100 0 0.83915 0.83915 0.83915 0.83915 0.83915
500 0.85957 0.86128 0.85872 0.85957 0.85617
1000 0.86383 0.86553 0.86638 0.86128 0.86213
2000 0.87234 0.87234 0.87234 0.86213 0.87149
4000 0.8783 0.88085 0.88085 0.86213 0.8766
8000 0.88085 0.8834 0.8817 0.86298 0.88426
16000 0.88766 0.88681 0.89021 0.86553 0.89021
32000 0.89787 0.89362 0.89617 0.86894 0.89362
Table 14: Full Results on TripClick (torso).
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.86791 0.86791 0.86791 0.86791 0.86791
500 0.87421 0.85787 0.87533 0.89561 0.87832
1000 0.87297 0.85656 0.87007 0.89834 0.88468
2000 0.87564 0.86304 0.87542 0.90043 0.88611
4000 0.88167 0.87094 0.88168 0.90521 0.89146
8000 0.88552 0.87471 0.88762 0.90846 0.89684
16000 0.8915 0.87811 0.89261 0.911 0.90002
32000 0.892 0.87517 0.89319 0.91485 0.90247
Recall@10 0 0.93622 0.93622 0.93622 0.93622 0.93622
500 0.94158 0.93736 0.94238 0.94385 0.94226
1000 0.94305 0.93727 0.94163 0.94312 0.94523
2000 0.94305 0.9397 0.9424 0.94194 0.94457
4000 0.94484 0.9412 0.9444 0.9424 0.9459
8000 0.94635 0.94284 0.94723 0.9434 0.94892
16000 0.94806 0.94366 0.9481 0.94293 0.95087
32000 0.9486 0.94342 0.94868 0.93992 0.94965
Success@10 0 0.97735 0.97735 0.97735 0.97735 0.97735
500 0.98305 0.97705 0.9835 0.98755 0.98455
1000 0.98305 0.9772 0.98185 0.988 0.9868
2000 0.9829 0.979 0.9826 0.98785 0.98545
4000 0.9847 0.9799 0.9838 0.98875 0.9865
8000 0.985 0.9805 0.9859 0.98965 0.9883
16000 0.98515 0.9802 0.9853 0.98935 0.9892
32000 0.9838 0.9799 0.98455 0.988 0.98695
NDCG@20 0 0.8723 0.8723 0.8723 0.8723 0.8723
500 0.87815 0.86211 0.87925 0.89912 0.88211
1000 0.87689 0.86165 0.87416 0.90187 0.88805
2000 0.88005 0.86781 0.87991 0.90409 0.89001
4000 0.88609 0.87588 0.88641 0.90866 0.89527
8000 0.88992 0.87945 0.89184 0.91165 0.90086
16000 0.89561 0.88274 0.89676 0.91402 0.90385
32000 0.89564 0.87916 0.89703 0.91796 0.90634
Recall@20 0 0.95048 0.95048 0.95048 0.95048 0.95048
500 0.95453 0.95144 0.95523 0.95469 0.95459
1000 0.95569 0.95407 0.95477 0.95399 0.95601
2000 0.95721 0.95493 0.95659 0.95319 0.95708
4000 0.95903 0.95695 0.9596 0.95303 0.95791
8000 0.96049 0.95829 0.96069 0.95324 0.96156
16000 0.96127 0.95902 0.96142 0.95232 0.96278
32000 0.96026 0.95607 0.96086 0.94967 0.96195
Success@20 0 0.9859 0.9859 0.9859 0.9859 0.9859
500 0.9901 0.98635 0.9904 0.9925 0.9904
1000 0.9898 0.98755 0.9895 0.99265 0.99175
2000 0.9907 0.98755 0.9898 0.9928 0.99175
4000 0.99145 0.98845 0.99205 0.9934 0.99205
8000 0.9925 0.9892 0.9922 0.9937 0.99355
16000 0.9919 0.9895 0.99175 0.99355 0.99355
32000 0.98965 0.98605 0.9901 0.99295 0.99235
NDCG@100 0 0.87636 0.87636 0.87636 0.87636 0.87636
500 0.88261 0.86725 0.88371 0.90305 0.88639
1000 0.88132 0.8665 0.87887 0.90575 0.89229
2000 0.88435 0.87259 0.88431 0.90809 0.89427
4000 0.89035 0.88029 0.89042 0.91256 0.89958
8000 0.89399 0.88387 0.89579 0.91537 0.90469
16000 0.89971 0.88716 0.90076 0.91775 0.90758
32000 0.89989 0.88413 0.90125 0.92182 0.91042
Recall@100 0 0.96671 0.96671 0.96671 0.96671 0.96671
500 0.97215 0.97231 0.97283 0.97047 0.9714
1000 0.97325 0.97381 0.97341 0.96951 0.97267
2000 0.97417 0.97459 0.97396 0.96964 0.97374
4000 0.97595 0.97473 0.97535 0.96888 0.97511
8000 0.97648 0.97609 0.97618 0.96812 0.97659
16000 0.97727 0.97661 0.97698 0.96719 0.97709
32000 0.97751 0.9762 0.97807 0.96518 0.97812
Success@100 0 0.99205 0.99205 0.99205 0.99205 0.99205
500 0.99565 0.9943 0.9958 0.99625 0.99565
1000 0.9955 0.9949 0.99565 0.99655 0.9958
2000 0.99505 0.99475 0.99505 0.99745 0.9958
4000 0.9955 0.99445 0.9949 0.99745 0.9964
8000 0.9958 0.99505 0.9955 0.99685 0.99655
16000 0.9958 0.9949 0.9958 0.9967 0.9964
32000 0.9958 0.99505 0.9961 0.9967 0.99685
Table 15: Full Results on FEVER.
Metric #iterations FT RP25 InfoBatch SP25 DP
NDCG@10 0 0.78999 0.78999 0.78999 0.78999 0.78999
500 0.79112 0.78914 0.78982 0.79181 0.79469
1000 0.79348 0.7896 0.79108 0.79162 0.79549
2000 0.79611 0.79243 0.79308 0.79228 0.79771
4000 0.79654 0.79531 0.79701 0.79543 0.80125
8000 0.80141 0.79758 0.80054 0.79639 0.80721
16000 0.80386 0.79894 0.80298 0.80032 0.81139
32000 0.80585 0.80083 0.80647 0.80275 0.81208
Recall@10 0 0.76583 0.76583 0.76583 0.76583 0.76583
500 0.77434 0.77286 0.77286 0.74477 0.77056
1000 0.77576 0.77441 0.77596 0.73943 0.774
2000 0.77981 0.77873 0.7788 0.74099 0.77731
4000 0.78379 0.7815 0.78325 0.74105 0.78325
8000 0.78899 0.78379 0.78778 0.74038 0.78771
16000 0.79433 0.78656 0.79318 0.74207 0.79365
32000 0.80088 0.78623 0.80068 0.74362 0.79953
Success@10 0 0.95179 0.95179 0.95179 0.95179 0.95179
500 0.95935 0.95881 0.95895 0.95422 0.96003
1000 0.96057 0.95935 0.95989 0.95409 0.96057
2000 0.96259 0.96003 0.96084 0.95409 0.96138
4000 0.96138 0.95868 0.9603 0.95395 0.96502
8000 0.96111 0.95841 0.96016 0.95449 0.96597
16000 0.96003 0.95787 0.96003 0.95463 0.96637
32000 0.96124 0.9576 0.96084 0.95544 0.96529
NDCG@20 0 0.80366 0.80366 0.80366 0.80366 0.80366
500 0.80525 0.80333 0.80421 0.80514 0.80885
1000 0.808 0.80409 0.80541 0.80501 0.80945
2000 0.8101 0.80683 0.80703 0.80565 0.81199
4000 0.81002 0.80926 0.81084 0.80884 0.81438
8000 0.81462 0.81148 0.81433 0.80939 0.82042
16000 0.81756 0.81293 0.81682 0.81337 0.82424
32000 0.81933 0.81502 0.81977 0.81546 0.82478
Recall@20 0 0.80743 0.80743 0.80743 0.80743 0.80743
500 0.81695 0.8154 0.81614 0.78521 0.81303
1000 0.81938 0.81783 0.81924 0.78008 0.81594
2000 0.82228 0.82208 0.821 0.7819 0.82019
4000 0.82458 0.82316 0.82492 0.78217 0.82357
8000 0.82876 0.82606 0.82924 0.78015 0.82829
16000 0.83565 0.82917 0.83477 0.78143 0.83268
32000 0.84159 0.82937 0.84126 0.78204 0.83808
Success@20 0 0.96705 0.96705 0.96705 0.96705 0.96705
500 0.97434 0.97407 0.97394 0.96772 0.97556
1000 0.97596 0.97529 0.97434 0.96718 0.97556
2000 0.97623 0.97515 0.97448 0.96678 0.97637
4000 0.97556 0.97448 0.97556 0.96691 0.97623
8000 0.97529 0.97218 0.97583 0.96691 0.97826
16000 0.97637 0.97259 0.97569 0.9684 0.97893
32000 0.97691 0.97205 0.97569 0.96907 0.97812
NDCG@100 0 0.82198 0.82198 0.82198 0.82198 0.82198
500 0.82227 0.82079 0.82131 0.82394 0.82592
1000 0.82477 0.8212 0.82234 0.82389 0.82651
2000 0.8265 0.82336 0.82383 0.82444 0.82881
4000 0.82664 0.82613 0.82728 0.82724 0.83099
8000 0.83127 0.82852 0.83067 0.828 0.83651
16000 0.8334 0.82963 0.83301 0.83107 0.84023
32000 0.83481 0.83163 0.83551 0.8332 0.84056
Recall@100 0 0.88292 0.88292 0.88292 0.88292 0.88292
500 0.88785 0.88805 0.88717 0.86313 0.88386
1000 0.8892 0.88926 0.89001 0.8582 0.88724
2000 0.89061 0.89129 0.89061 0.85976 0.89041
4000 0.89318 0.89332 0.89352 0.85814 0.89271
8000 0.89757 0.89608 0.89683 0.85658 0.89534
16000 0.90095 0.89797 0.90149 0.85442 0.89878
32000 0.90554 0.89764 0.90628 0.85571 0.90365
Success@100 0 0.98825 0.98825 0.98825 0.98825 0.98825
500 0.99082 0.99095 0.99068 0.98825 0.99122
1000 0.99149 0.99122 0.99055 0.98798 0.99109
2000 0.99149 0.99109 0.99176 0.98785 0.9919
4000 0.99163 0.99149 0.99176 0.98866 0.99203
8000 0.9919 0.99217 0.99082 0.98866 0.99203
16000 0.99203 0.9919 0.99203 0.98866 0.99298
32000 0.99257 0.99109 0.99203 0.98879 0.9923
Table 16: Full Results on HotpotQA.
Table 17: Qwen3-Embedding-0.6B on NFCorpus.
Metric Pretrained FT InfoBatch SP DP
MRR@1 0.458 0.502 0.511 0.533 0.517
MRR@5 0.543 0.572 0.580 0.595 0.587
MRR@10 0.550 0.581 0.587 0.605 0.595
MRR@20 0.555 0.584 0.590 0.607 0.598
MRR@50 0.556 0.586 0.591 0.609 0.600
MRR@100 0.557 0.586 0.592 0.609 0.600
NDCG@5 0.440 0.479 0.480 0.492 0.486
NDCG@10 0.441 0.479 0.478 0.487 0.479
NDCG@20 0.461 0.489 0.484 0.489 0.492
NDCG@50 0.517 0.539 0.540 0.545 0.541
NDCG@100 0.577 0.617 0.617 0.615 0.618
Recall@1 0.061 0.065 0.068 0.067 0.065
Recall@5 0.136 0.162 0.165 0.155 0.169
Recall@10 0.171 0.236 0.237 0.214 0.234
Recall@20 0.211 0.314 0.308 0.265 0.311
Recall@50 0.267 0.416 0.417 0.357 0.408
Recall@100 0.329 0.518 0.517 0.443 0.502
Success@1 0.458 0.502 0.511 0.533 0.517
Success@5 0.672 0.690 0.697 0.693 0.697
Success@10 0.721 0.755 0.749 0.768 0.762
Success@20 0.793 0.796 0.796 0.793 0.805
Success@50 0.836 0.842 0.845 0.851 0.864
Success@100 0.889 0.889 0.885 0.885 0.892
Table 18: Qwen3-Embedding-0.6B on FiQA. The pretrained model outperforms all finetuned methods, likely due to high-quality financial data in its pretraining corpus.
Metric Pretrained FT InfoBatch SP DP
MRR@1 0.469 0.392 0.394 0.390 0.384
MRR@5 0.543 0.465 0.466 0.464 0.464
MRR@10 0.555 0.474 0.473 0.473 0.471
MRR@20 0.559 0.479 0.478 0.477 0.476
MRR@50 0.561 0.480 0.480 0.480 0.478
MRR@100 0.561 0.481 0.481 0.480 0.479
NDCG@5 0.469 0.420 0.428 0.423 0.423
NDCG@10 0.511 0.454 0.455 0.456 0.450
NDCG@20 0.540 0.483 0.484 0.478 0.476
NDCG@50 0.568 0.506 0.508 0.506 0.501
NDCG@100 0.587 0.517 0.520 0.522 0.515
Recall@1 0.245 0.199 0.197 0.205 0.192
Recall@5 0.453 0.377 0.384 0.377 0.377
Recall@10 0.552 0.446 0.439 0.448 0.432
Recall@20 0.633 0.510 0.508 0.504 0.495
Recall@50 0.724 0.574 0.583 0.584 0.565
Recall@100 0.798 0.612 0.623 0.634 0.611
Success@1 0.469 0.392 0.394 0.390 0.384
Success@5 0.674 0.586 0.596 0.590 0.583
Success@10 0.761 0.657 0.644 0.656 0.634
Success@20 0.821 0.715 0.728 0.713 0.707
Success@50 0.886 0.767 0.781 0.781 0.769
Success@100 0.927 0.802 0.816 0.832 0.802
Table 19: Qwen3-Embedding-0.6B on ANTIQUE.
Metric Pretrained FT InfoBatch SP DP
MRR@1 0.615 0.580 0.625 0.640 0.615
MRR@5 0.718 0.682 0.707 0.738 0.696
MRR@10 0.724 0.689 0.716 0.746 0.703
MRR@20 0.727 0.692 0.718 0.747 0.706
MRR@50 0.728 0.693 0.719 0.748 0.707
MRR@100 0.728 0.693 0.719 0.748 0.707
NDCG@5 0.519 0.496 0.502 0.546 0.510
NDCG@10 0.518 0.496 0.506 0.540 0.520
NDCG@20 0.570 0.546 0.549 0.582 0.567
NDCG@50 0.646 0.617 0.624 0.658 0.636
NDCG@100 0.690 0.663 0.665 0.707 0.674
Recall@1 0.072 0.071 0.069 0.074 0.074
Recall@5 0.214 0.202 0.197 0.231 0.207
Recall@10 0.304 0.270 0.264 0.317 0.284
Recall@20 0.398 0.340 0.329 0.396 0.353
Recall@50 0.510 0.444 0.431 0.514 0.450
Recall@100 0.585 0.516 0.495 0.597 0.513
Success@1 0.615 0.580 0.625 0.640 0.615
Success@5 0.855 0.835 0.840 0.885 0.820
Success@10 0.905 0.890 0.905 0.935 0.870
Success@20 0.940 0.930 0.930 0.950 0.915
Success@50 0.980 0.955 0.950 0.975 0.950
Success@100 0.980 0.965 0.955 0.980 0.955
Table 20: Qwen3-Embedding-0.6B on TriviaQA.
Metric Pretrained FT InfoBatch SP DP
MRR@1 0.519 0.519 0.517 0.538 0.547
MRR@5 0.602 0.615 0.611 0.620 0.634
MRR@10 0.611 0.624 0.619 0.627 0.642
MRR@20 0.614 0.627 0.622 0.630 0.645
MRR@50 0.616 0.628 0.624 0.631 0.646
MRR@100 0.616 0.628 0.624 0.631 0.647
NDCG@5 0.461 0.483 0.478 0.489 0.501
NDCG@10 0.467 0.489 0.484 0.493 0.504
NDCG@20 0.502 0.522 0.515 0.525 0.535
NDCG@50 0.564 0.587 0.582 0.584 0.598
NDCG@100 0.605 0.632 0.626 0.623 0.640
Recall@1 0.086 0.094 0.093 0.093 0.097
Recall@5 0.236 0.267 0.263 0.253 0.271
Recall@10 0.316 0.360 0.354 0.334 0.363
Recall@20 0.403 0.462 0.453 0.421 0.461
Recall@50 0.514 0.590 0.584 0.529 0.588
Recall@100 0.591 0.677 0.669 0.603 0.669
Success@1 0.519 0.519 0.517 0.538 0.547
Success@5 0.735 0.766 0.759 0.752 0.766
Success@10 0.795 0.826 0.821 0.803 0.826
Success@20 0.844 0.871 0.864 0.848 0.867
Success@50 0.886 0.912 0.905 0.888 0.912
Success@100 0.911 0.935 0.927 0.911 0.931